OpenStudy (anonymous):

This may end up being stupid but Solving equations by Using the Zero Product Rule I did it this way is this OK? Or...? problem is : x(x+4)(10x-3)=0 x+4-4=0-4 10x-3+3=0+3 10/10x=3/10 x=3/10 then x^2+4x+10x^2-3x=0 set all x's to zero? Or is that entirely the wrong way to do this? Cos I am not sure what I should be doing. Thanks.

Nnesha (nnesha):

first you don't have subtract 4 yet x(x+4)(10x-3)=0 its mean all is equal to 0 like this x=0 (x+4)=0 10x -3 = 0 now solve for x

OpenStudy (cwrw238):

x(x+4)(10x-3)= 0 by the Zero Produce rule either one of the factors has to be zero - there are 3 factors x , x+4 and 10x - 3 so x = 0 or x+4 = 0 or 10x - 3 = 0 so solving these 3 equations ;- the solution set is x = 0 , -4 , 3/10

OpenStudy (anonymous):

Now, what do I do if it has a number where that x is? LIke, 5(y-.4)(y+2.1)=0 ?

OpenStudy (anonymous):

Is it the same thing?

OpenStudy (cwrw238):

no - its 5 times 9 y - 0.4) so y- 0.4 = 0 so y = 0.4

OpenStudy (cwrw238):

** thats not 9 thats (y - 0.4)